OpenStudy (anonymous):

convert the equation to this form: y=a(x-h)^2+k y=-6x^2+24x-21 PROBLEM

OpenStudy (anonymous):

y=-6(x-2)^2+3

OpenStudy (anonymous):

(x-2)^2= (x^2-4x+4)...*-6=-6x^2+24x-24...+3= -6x^2+24x-21... unless I majorly messed something up.

OpenStudy (anonymous):

THANKS! :)!!